answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

In a twenty question true false examination a student must achieve eight correct answers to pass if student answers randomly what is the probability that student will fail?

The probability this student will fail is the same as the probability that some other student will flip a fair coin 20 times and get less than 8 heads, i.e., more than 12 tails.

There are 2^20 possible different-looking sequences of 20 coinflips, which we assume all have equal probability.

Of those sequences, 1 has no heads at all, 20 have exactly 1 head, 190 have exactly 2 heads, ... and 77520 have exactly 7 heads.

So we sum up all those possible ways to fail and we get ...

... ...

I'm assuming that the student answers randomly, flipping a fair (50:50) coin on each question to choose "true" or "false".

In that special case, it doesn't matter how many of those twenty questions are true or how many are false.

(If the student answers randomly by flipping an unfair coin, say a 25:75 coin, then it does matter how many of those questions are true -- I'll let you figure that one out).

Why is only one normal distribution table needed to find any probability under the normal curve?

The normal distribution is symmetric about its mean; it increases from an asymptote with the x-axis below the mean until the mean whereupon it decreases until another asymptote with the x-axis the same distance above the mean. (This is not a linear increase/decrease, but a "bell" shape.)

As the distribution is symmetric about its mean, only tables up to the mean need be calculated/given in a table. The area under the curve between any two points can then be calculated.

For a normal distribution with mean µ and standard deviation σ, a z value is calculated for a given point x:

z = (x - µ) / σ

This z value is then used to look up the area in the given "half" tables, giving the area (probability) of the value lying between the mean and the given z value. If negative, z is below the mean, but for the table, the sign is ignored.

This can be expressed as:

area = normal(|z|)

where normal(z) is the value in the normal table at the given (positive) z value.

To calculate the area between two points (ie the probability that a value lies between two given values), their corresponding z values (z₁ and z₂) are first calculated and then combined viz:

  • If they are both on the same side of the mean (ie z₁ and z₂ have the same sign) then the area is given by:
area = | normal(|z₁|) - normal(|z₂|) |

  • If they are on opposite sides of the mean (ie z₁ and z₂ have different signs) then the area is given by:
area = normal(|z₁|) + normal(|z₂|)

Almost all of the normal distribution lies between ±4 standard deviations of the mean.

What are the 6 best numbers to use to win any lottery?

There is no set of "best numbers". If there were someone would have used them again and again to amass a fortune.

When tossing 5 coins simultaneously find the probability that at least 1 head is showing.?

This is easiest to solve by working out the probability that no heads show and subtracting this from 1 to give the probability that at least one head shows:

Assuming unbiased coins which won't land and stay on their edge, the probability of head = probability of tail = ½
→ probability no heads = probability 5 tails = ½^5 = 1/32
→ probability of at least one head = 1 - 1/32 = 31/32 = 0.96875 = 96.875 % = 96 7/8 %

What is the probability of 0?

The answer depends on what the experiment is.

What is the probability of rolling a number higher than 4?

There are six faces, and only three faces with a number higher than a four. So it is 3 chances in 6.

A bag contains 5 balls two balls are drawn at a random and found to be red what is the probability that all balls are red?


The probability that all balls are red is 0.50 or 50%.

EXPLANATION

Suppose there are 4 boxes with 5 balls each.

Box A has 2 red balls.

Box B has 3 red balls.

Box C has 4 red balls.

Box D has 5 red balls.

The probability of drawing at random 2 red balls for each box is:

Box A; P(2 red balls) = (2/5)∙(1/4) = 2/20 = 1/10

Box B; P(2 red balls) = (3/5)∙(2/4) = 6/20 = 3/10

Box C; P(2 red balls) = (4/5)∙(3/4) = 12/20 = 6/10

Box D; P(2 red balls) = (5/5)∙(4/4) = 20/20 = 10/10

Now, suppose we do the drawing of 2 balls experiment 10 times on each box

giving a total of 40 experiments. The probabilities calculated above are the

"expected" results, that is; out of the 40 experiments (drawing of 2 balls), two

red balls resulted:

1 time came from box A

3 times from box B

6 times from box C

10 times from box D

Notice that from the 40 experiments, 20 result in 2 red balls.

From here we have that when drawing 2 red balls the probability that it came

from a box containing 2 red balls (box A), 3 red balls (box B), 4 red balls (box C)

or 5 red balls is:

P(A) = 1/20 = 0.05 = 5%

P(B) = 3/20 = 0.15 = 15%

P(C) = 6/20 = 0.30 = 30%

P(D) = 10/20 = 0.50 = 50%

What is an example of independence events?

An example of independent events is rolling a die and flipping a coin. The roll of the die has no affect on the flipping of the coin and the converse is true also. Another example is the rolling of a die followed by the rolling of a die - they are sequentially unrelated - so they are independent.

A group's reported are more likely to become equal to the statistical probability of loss?

the larger the group, the more likely the statistical probability of loss will be equal

The probability that a tennis set will go to a tie breaker is 10 percent what is the probability that two of the three sets will go to a tie breaker?

The outcomes for each set is not independent. For a set to go into a tie breaker the players need to be very evenly matched. If they are, there is a greater probability that the sets will go into tie breaker. If you disregard that aspect of realitythen the answer is 0.0270, approx.

What is the difference between experimental and theoretical physics?

When theoretical physicists work on equations and don't test their hypothesis, experimental physicists test their hypothesis and verify their conclusion. Usually theoretical physicists work on things like black holes and string-theory when experimental physicists work on Newtonian laws.

What are 5 things that are matter and five things that are not matter?

Five things that aren't matter are: sound, light, heat, gravity, and speed. Five things that are matter are air, water, foods, pencils, and paper.

How many ways are there to rearrange the letters in marmalade?

Calculate 9! and put that into the numerator. In the denominator, you need to account for repeated letters; since there are 2 "m" and 3 "a", the denominator should be 2! times 3!.

Note: The exclamation refers to the factorial.

How are ratio and probability the same?

A ratio is a comparison of the relative size of two different things. Probability is the change that something will (or will not) occur.

Probability can be expressed as a ratio of Yes to No (or, "will occur" to "won't occur"). That is, Probability is the relative size of Yes to No. So, if something is said to have a 60% Probability of occurring, what that is indicating is that, out of 100 tries, 60 will be the outcome indicated. While probability is usually expressed as a percentage, it is entirely possible to express it as a ratio. In the aforementioned example, a 60% Probability of occurrence could also be said to be a 60:40 (or, reduced, 3:2) ratio in favor of happening.

Trending Questions
Drew gave 12 marbles to Marcus each boy then had 25 marbles how many more marbles did drew have than Marcus at first? What are the rules of probability? Is chi-square test examines extent of errors between variables? In six decks of playing cards what is the probability of six seven and eight of the same suit to come up? What is the function of sartorius? What is the probability of getting a multiple of 3 or 4 in throwing a dice? What is the probability of getting heads when you toss a coin? Carmen flips two nickels at the same time. Is the probability of getting two heads the same as that of getting one head and one tail? What is the probability that a state does not begin with a vowel? If you roll a 6 faced dice how may outcomes are there? A number cube is rolled 2 times in a row What is the probability of rolling a 3 both times? Tthe probability that two heterozygous parents for widow's peak will have a child with a straight hairline is what percent? What are the probability genotype of the parents? How many combinations of two numbers are there in 16 numbers? What is the probability of getting a 2 or 12 on two dice? A card is drawn at random from a deck of cards Find the probability of getting the King of heart? Why is the event described in the article taught in schools today What are the lasting lessons that we can learn from studying this event? Two coins are tossed What is the probability of both coins are heads? If two coins are tossed what is the possivle outcome? How many ace are in a deck?